OpenStudy (anonymous):

i understand that The quantity x squared plus 2 times x minus 3 all over the quantity x squared minus 1, end quantity = the quantity x plus 3, end quantity, times the quantity x minus 1, end quantity, all over the quantity x squared minus 1. has no GCF, but how in the heck do you get that as the simplified version

OpenStudy (agent0smith):

Don't write an equation like that and expect us to want to decipher it. Use brackets and such.

OpenStudy (anonymous):

ok sorry i just copy and pasted it, it came out like tht x^2+2x-3/x^2-1

OpenStudy (anonymous):

thts the original equation, this is what it looks like simplified, (x+3)(x-1)/x^2-1

OpenStudy (jdoe0001):

\(\bf \cfrac{x^2+2x-3}{x^2-1}=\cfrac{(x+3)(x-1)}{x^2-1}\\ \quad \\ \textit{recall that }\qquad \color{blue}{a^2-b^2 = (a-b)(a+b)}\qquad thus\\ \quad \\ \cfrac{(x+3)(x-1)}{x^2-1}\implies \cfrac{(x+3)(x-1)}{x^2-1^2}\implies \cfrac{(x+3)(x-1)}{(x+1)(x-1)}\)

OpenStudy (anonymous):

i know how to get the bottom, but how do you get the top like that, do you divide the two out and thts what makes the one?

OpenStudy (jdoe0001):

hmmm how does the trinomial gets factored to 2 binomials?

OpenStudy (anonymous):

i see what happens to the x^2 and the -3 but not the 2x, brb i gotta get my dogs in

OpenStudy (jdoe0001):

\(\bf (x+3)(x-1)\implies x^2-x+3x-3\implies x^2-2x-3\)

OpenStudy (jdoe0001):

ahem... rather \(\bf (x+3)(x-1)\implies x^2-x+3x-3\implies x^2+2x-3\)

OpenStudy (anonymous):

so the x^2+2x-3 equals (x+3)(x-1) for no reason thts just how it is?

OpenStudy (jdoe0001):

well... anything wrong with the multiplication above?

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=WipYqupv9wY

OpenStudy (anonymous):

i'm not saying you're doing anything wrong, i just don't understand how you get those numbers thts all

OpenStudy (anonymous):

well thanks for helping but i'm gonna use kahn academy

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

i figured it out, i forgot to do the foil thing lol i'm dumb